The woods by Vince’s house have a direct path to his school. If he walks from his house through the woods, he walks approximately 26.9 meters to get to his school. Otherwise, from his house, he would need to walk south to the crosswalk then walk west to arrive at his school. If Vince walks the same distance south as he does west, how many meters south does Vince need to walk? Round your answer to the nearest tenth
Found 2 solutions by Boreal, stanbon:
Answer by Boreal(15235)

This is a 45-45-90 right triangle with the hypotenuse 26.9 meters. These triangles have each side 1-1 and hypotenuse sqrt (1+1) or sqrt (2)
each side is hypotenuse/sqrt(2) meters or 19.02 or 19.0 meters

Answer by stanbon(75887)

The woods by Vince’s house have a direct path to his school. If he walks from his house through the woods, he walks approximately 26.9 meters to get to his school. Otherwise, from his house, he would need to walk south to the crosswalk then walk west to arrive at his school. If Vince walks the same distance south as he does west, how many meters south does Vince need to walk? Round your answer to the nearest tenth
----
Sketch the picture.
You have a right triangle with sou361.th dist = x = west distance
and the hypotenuse is 26.9 meters
----
Equations:
x^2 + x^2 = 26.9^2
----
x^2 = 26.9^2/2 = 361.805
----
x = 19.02 meters
-----------
Cheers,
Stan H.
---------------